Elasticity of apple supply s                                          3

Own price elasticity of apple  demand        -2.5

Elasticity of pear supply                                  5

Own Price elasticity of [pear demand             -3

Analyze the percentage changes in prices and quantity in the following scenarios

1.     A rise in the price of oranges causes the demand of apples to increase by 15%




1
Expert's answer
2021-04-27T19:05:54-0400

Assuming that Oranges and Apples are substitute goods, a rise in the price of oranges causes the demand for apples to increase. In Equilibrium Displacement Model (EDM) only partial equilibrium is considered where a shock affects only one market. However, this is always not the case especially when the goods are complementary or substitute to each other.



Since the demand is going through a change, we will use the following formula to calculate the percentage change.


Price elasticity of apple demanded "=\frac{percentage\ change\ in\ quantity\ demanded}{percentage\ change\ in\ price\ for\ oranges}"


The price elasticity of apple demanded= -2.5


Percentage change in quantity demanded= 15%


Therefore,


"-2.5= \frac{15\%}{\%change\ in\ price\ for\ oranges}"


"\%change\ in\ price\ for\ oranges\ =\frac{15\%}{2.5}"


"=6\%"

The price of Oranges will, therefore, increase by 6% from its current price. On the other hand, the demand for Oranges will start to decrease with an increase in price.


Explanation.


In this case, a rise in demand for Apples will cause a feedback effect on the Orange market.

Therefore, higher Apple demand will lead to an increase in Apple prices which will cause some consumers to demand more oranges. This will make the prices of oranges increase further because more consumers now prefer buying Oranges instead of Apples. Due to an increase in price, the demand for Oranges will start to decrease. An increase in price for Oranges will lead to a decrease in the quantity demanded.